# Super-Resolved Canopy Height Mapping from Sentinel-2 Time Series Using LiDAR HD Reference Data across Metropolitan France

### v1

@@ 4.2.2 Qualitative Results
4.2.3 Temporal monitoring
((a)) 10m
@@ 4.2.3 Temporal monitoring
((c)) 2.5m
((d)) SPOT 2022
((e)) SPOT 2023
((f)) ALS change map at 1.5m
Figure 8: Change detection example centered on 49.187°N and 2.573°E. Subfigures (a)–(c) present the prediction results of our model at 10 m, 5 m, and 2.5 m resolutions, respectively. Subfigure (d) and (e) show reference VHR SPOT images acquired corresponding to 2022 and 2023 acquisition years, while (f) depicts the reference ALS-derived height change map between September 2023 and February 2022 at 1.5 m resolution. Both SPOT and ALS images are taken from Open-Canopy repository (Fogel et al., 2025).
Figure 8 illustrates fine-scale change detection in the Chantilly forest. Although our model was not explicitly trained to predict non-vegetation areas – assigning them a default height of zero – it still produces values close to zero in these regions, making change detection feasible. Despite this constraint, the loss of vegetation is clearly captured, particularly at the 2.5 m and 5 m resolution. The reference VHR SPOT images from 2022 and 2023, as well as the ALS-derived height change map at 1.5 m resolution (between February 2022 and September 2023), are taken from the Open-Canopy repository (Fogel et al., 2025).
((a)) BD ORTHO 2018
((b)) BD ORTHO 2021
((c)) BD ORTHO 2024
((d)) Pinus Pinaster regrowth profile.
Figure 9: Regrowth monitoring of Pinus Pinaster example centered on 44.894°N and 1.13°W. Subfigures (a)–(c) present the BD ORTHO images for 2018, 2021 and 2024 year. Subfigure (d) shows the regrowth profile for our model at 10 m, 5 m, and 2.5 m resolutions. The thick line corresponds to a median value within the delineated parcel, with the margins extending to $25^{th}$ and $75^{th}$ percentiles.
Figure 9 illustrates the regrowth profile of Pinus pinaster in the Landes of Gascogne. In this region, sanitary clear-cuts are regularly performed, followed by the planting of young trees, making the monitoring of early-stage growth particularly important. We delimitate a forest parcel with young pine trees to follow its regrowth profile.
The plot shows the median tree height per parcel, accompanied by the 25th and 75th percentiles of parcel‐level height values. The 2.5 m and 5 m predictions exhibit very similar behaviour across all ages. In contrast, the 10 m model reports higher height values for young stands. This discrepancy is not necessarily an error: at finer resolutions (2.5 m and 5 m), young trees are sparsely distributed and gaps between them are visible, whereas at 10 m these gaps disappear at 10 m resolution, since the model uses the 95th percentile of point heights within each pixel, yielding higher height estimates. As stands grow older, this difference gradually diminishes. The similarity between the 2.5 m and 5 m curves for young stands suggests that the 2.5 m model captures only limited additional structural detail compared to the 5 m model.
Although the predicted growth curve shows occasional small drops, the overall tendency remains within the expected error range. These slight decreases may reflect periods of drought or other meteorological events that affect the spectral response of the vegetation, leading to temporarily lower predicted heights.
Finally, the predicted regrowth curve aligns with the expected growth dynamics of young Pinus pinaster in the Landes forest (Chevalier et al., 2024), which reaches approximately 1 m per year, depending on the climatic conditions. This confirms that our method is well suited for monitoring early-stage forest regrowth.
Moreover, we evaluated the effectiveness of temporal conditioning by explicitly generating predictions at different dates throughout the year. Although no reference data are available at a monthly frequency, the expected seasonal growth pattern can still be observed when comparing predictions across the specified dates.
5 Conclusion
In this paper, we have introduced THREASURE-Net, a novel end-to-end framework for super-resolved tree canopy height estimation. Our approach generates high-resolution annual predictions at 2.5 m and 5 m directly from Sentinel-2 time series with 10 m resolution, supervised with LiDAR HD–derived canopy height data. THREASURE-Net outperforms existing Sentinel-2–based methods for forest height prediction and achieves performances comparable to approaches relying on 1.5 m SPOT-6/7 imagery. Our model achieves mean absolute errors of 2.62 m, 2.72 m, and 2.88 m at 2.5 m, 5 m, and 10 m resolution, respectively, and can be used for change detection and tree growth monitoring.
Finally, we plan to integrate a tree/non-tree classification module, enabling a fully end-to-end system that does not rely on an external forest mask.
Acknowledgements
Ekaterina Kalinicheva, Florian Mouret and Milena Planells were supported by the PC ALAMOD from the PEPR FairCarbon. Florian Helen and Stéphane Mermoz were supported by a “R&T: Estimation annuelle de paramètres bio-géophysiques des forêts françaises et leurs changements par IA” grant from the Centre National d’Études Spatiales (CNES) (Marché n°5700013599).
We would like to express our sincere gratitude to Julien Michel and Jordi Inglada for their assistance and valuable advice on deep learning methods, without which this work would not have been possible.

### v2

@@ 4.2.2 Qualitative Results
4.3 Ablation Study
To justify the benefits of the selected loss functions, we assess their individual contributions relative to a baseline L1 loss. We conduct an ablation study on the 5 m resolution model considering four variants: (i) training with the standard L1 (MAE) loss only; (ii) training with the patch-average L1 loss only (here referred to as patch-weighted MAE, pwMAE); (iii) the proposed configuration combining wGDL with pwMAE; and (iv) the proposed configuration combining wGDL with pwMAE using an increased non-tree pixel weight of $w_{\text{non-tree}}=0.5$. Note that in both models (i) and (ii) we apply a non-tree weight of $w_{\text{non-tree}}=0.1$, as in our proposed configuration (iii).
Table 3: Ablation study of different loss functions.
Model Variant MAE rMSE MAE (%) $R^{2}$ IoU
MAE 2.76 3.74 26.2 0.71 85.6
pwMAE 2.76 3.73 25.9 0.71 86.4
wGDL + pwMAE (Our) 2.70 3.66 25.3 0.72 86.7
wGDL + pwMAE $w_{\text{non-tree}}=0.5$ 2.90 3.93 27.4 0.67 87.3
Figure 8: FAP profile for the target image, SR image produced by our model at 5 m resolution with the chosen hyperparameters and two variation of our model: the one using only MAE loss, for training, and the other using only pwMAE for training, without wGDL loss.
The quantitative results are presented in Table 3. We observe a slight improvement of pwMAE over MAE. Moreover, the improvement of our full model compared to the configuration without the wGDL loss is non-negligible. Finally, increasing the weight of non-tree pixels to 0.5 improves the classification mask but reduces regression performance.
Figure 8 presents FAP profiles for configurations (i), (ii), and (iii), where we observe improved frequency reconstruction when using the wGDL loss compared to the variants without it. Those observations prove the efficacy of the proposed model configuration.
4.4 Comparison with other methods using independent LiDAR reference map
We evaluate our model at different spatial resolutions and compare it with three recent canopy height mapping approaches using an independent ALS dataset acquired in 2020. Specifically, we compare against FORMS-T (Schwartz et al., 2025) at 10 m resolution, Open-Canopy (Fogel et al., 2025) at 1.5 m resolution, and the Global Canopy Height (GCH) Model (Lang et al., 2023) at 10 m resolution.
The evaluation is conducted over the Landes region in southwestern France. For the competing methods, we use canopy height maps provided directly by the authors of the corresponding models. The reference LiDAR data cover an area of 580 ha.
Evaluation is performed at the native resolution of each canopy height product by reprojecting and aggregating the reference LiDAR height data to the corresponding grid.
((a)) Reference Height, 1 m
((b)) VHR Image
((c)) Our, 10 m
((d)) FORMS-T, 10 m
((e)) Our, 5 m
((f)) Lang, 10 m
((g)) Our, 2.5 m
((h)) Open-Canopy, 1.5 m
Figure 9: Tree height prediction of 580 ha area centred on 45.127°N and 0.901°W. Comparison of a) independent reference ALS-derived heights and our model at three resolutions—b) 10 m, c) 5 m, and d) 2.5 m—with concurrent approaches: d) FORMST-T (Schwartz et al., 2025), e) Global Canopy Height Model (Lang et al., 2023), and f) Open-Canopy (Fogel et al., 2025). Zoomed-in views (top left) highlight differences in fine details among predictions on the eastern side of the study area.
Table 4: Quantitative results for our models of 10, 5 and 2.5 m resolutions on the independent area of 580 ha in Landes region in South-West of France. We also report the performance of two other recent state-of-the-art algorithms: FORMS-T (Schwartz et al., 2025) at 10 m, Global Canopy Height Model (Lang et al., 2023) at 10 m, and Open-Canopy (Fogel et al., 2025) at 1.5 m resolution.
Model Res., m MAE, m MAE, % $R^{2}$ RMSE, m IoU, %
THREASURE-Net 10 2.26 35.4 0.74 2.98 97.6
5 2.32 57.8 0.70 3.22 94.5
2.5 2.45 79.2 0.65 3.49 92.5
FORMS-T 10 2.26 46.4 0.72 3.10 86.5
GCH 10 4.08 85.3 0.235 5.08 90.7
Open-Canopy 1.5 2.41 49.2 0.66 3.41 97.7
Figure 9 shows the qualitative results for the study area, while Table 4 summarizes the corresponding quantitative results. For our algorithm, for tree/non-tree classification, we directly use predicted binary mask while for other algorithms, the 2 m threshold was applied to the tree height prediction to isolate “non-tree” class. The Global Canopy Height model gave the worst results compared to other models, likely because it was trained with global data, contrary to the other algorithms designed specifically for the French temperate forests. This result suggests that region-specific models provide advantages over globally trained models for local-scale studies. From visual observation, it can be noted that Open-Canopy underestimates high values and flattens spatial textures, providing more homogeneous values within parcels, contrary to the other algorithms. This is reflected in the higher rMSE compared to FORMS-T and our models at 5 and 10 m, confirming that the Open-Canopy model tends to predict values close to the dataset mean, which results in poorer performance for extreme values, as noted earlier. Moreover, for 10 m resolution, our model visually gives much sharper textures than FORMS-T, which can possibly be explained by the use of wGDL loss.
High rMAE values are observed for our models at 5 and 2.5 m resolution, as well as for the concurrent approaches, despite their low absolute MAE. This can be explained by the prevalence of narrow gaps within the study area – often smaller than the pixel resolution – which are represented by low reference values and therefore increase the sensitivity of relative error to small absolute deviations, as the algorithms tend to smooth over these gaps at finer resolutions. As observed, our model at 10 m resolution yields the best rMSE, likely because the number of such gaps decreases at this coarser resolution.
Finally, Open-Canopy yields the best results for tree/non-tree classification, giving the highest IoU, which can be explained by the use of VHR input images that provide better contour delineation and detection of small gaps.
5 Conclusion
In this paper, we have introduced THREASURE-Net, a novel end-to-end framework for super-resolved tree canopy height estimation. Our approach generates high-resolution annual predictions at 2.5 m and 5 m directly from Sentinel-2 time series with 10 m resolution, supervised with LiDAR HD–derived canopy height data. THREASURE-Net outperforms existing Sentinel-2–based methods for forest height prediction and achieves performance comparable to approaches relying on 1.5 m SPOT-6/7 imagery. Our model achieves mean absolute errors of 2.63 m, 2.70 m, and 2.89 m at 2.5 m, 5 m, and 10 m resolution, respectively, and can be used for tree loss detection and tree growth monitoring.
Finally, we plan to improve a tree mask classification module, enabling model to better delimitate gaps.
Acknowledgements
Ekaterina Kalinicheva, Florian Mouret and Milena Planells were supported by the PC ALAMOD from the PEPR FairCarbon and received government funding managed by the Agence Nationale de la Recherche under the France 2030 program. Florian Helen and Stéphane Mermoz were supported by a “R&T: Estimation annuelle de paramètres bio-géophysiques des forêts françaises et leurs changements par IA” grant from the Centre National d’Études Spatiales (CNES) (Marché n°5700013599).
We would like to express our sincere gratitude to Julien Michel and Jordi Inglada for their assistance and valuable advice on deep learning methods, without which this work would not have been possible. We also gratefully acknowledge Jerôme Bock from ONF for providing the independent ALS-derived canopy height data used in this study for forest disturbance detection task.
Appendix A Sylvoécorégions of France
Figure 10: Sylvoécorégions of France, illustrating the ecological diversity of forest types across the country.
Appendix B Tree growth evaluation
We evaluate our algorithm on the tree regrowth analysis task. For this task, we choose an entire Sentinel - 2 tile 30TXQ corresponding to the Landes of Gascogne in the South-West of France, mostly populated by Pinus pinaster. First, we produce time series of tree height maps and associated binary tree classification masks with THREASURE-Net for 2018-2024 years for three spatial resolutions – 10, 5, and 2.5 m. We analyze tree classification masks in a temporal context to determine the test area by selecting pixels that were classified as tree in at least three timestamps.
Second, we apply the Mann–Kendall (MK) test Mann (1945) to the resulting time series. The Mann–Kendall test is a non-parametric statistical test used to detect monotonic trends in time series data. The test statistic varies between -1 (strong decrease) and +1 (strong increase). The MK test is applied at the parcel level using cadastral boundaries combined with the vegetation mask. Statistically insignificant results ($p_{value}\geq 0.05$) are discarded. The remaining values are grouped into three trend intervals: decreasing [1,0.5)[-1,-0.5), stable [0.5,0.5)[-0.5,0.5), and increasing $[0.5,1]$.
Within the tree classification mask, 38% of pixels are statistically insignificant. This proportion is expected given the relatively short length of the time series, which may be insufficient to capture annual growth trends, especially when the mean absolute error (MAE) exceeds the typical yearly regrowth signal. These pixels may correspond to disturbed areas, height estimates with higher uncertainty, or non-vegetated surfaces.
Furthermore, the Mann–Kendall index has limitations when interpreting forest disturbances occurring at the beginning or in the middle of the time series, as such events do not necessarily produce a monotonic trend. In practice, only disturbances occurring toward the end of the time series tend to be reflected by low index values.
For the remaining statistically significant pixels, the results are consistent across all spatial resolutions. Approximately 10% of pixels exhibit low Mann–Kendall values, while about 90% show high values, indicating increasing trends. In contrast, the proportion of medium trends [0.5,0.5)[-0.5,0.5) is extremely small (¡ 0.001%).
The near absence of medium trends suggests that almost no parcels exhibit a constant monotonic signal throughout the time series. This behavior is expected for forest vegetation, where growth dynamics are typically non-linear, and it also supports the quality of the tree classification mask.
Figure 11 shows the tree growth of parcels exhibiting high positive trends across the three spatial resolutions. Values are reported as the median, along with the 25th and 75th percentiles of statistically significant pixels with high trends. Interestingly, the years showing weaker or negative growth in the time series coincide with extreme climatic events, which may influence vegetation physiology and leaf composition, thereby affecting canopy reflectance properties that are directly related to height estimation. In particular, 2022 was marked by a severe drought in France, while 2024 was reported as one of the hottest summers recorded in the country since 1900. Although the variations observed in the predicted heights remain within the expected error margins of the model, these temporary decreases may still reflect the impact of extreme climatic conditions on vegetation growth. Moreover, while 10 m curve looks smoother, 2.5 and 5  curves exhibit identical behavior. Finally, the predicted growth curves align with the expected growth dynamics of young Pinus pinaster in the Landes forest (Chevalier et al., 2024), which reaches approximately 1 m per year, depending on the climatic conditions.
Consequently, deviations from the presented expected growth dynamics may reveal disturbance events, motivating the analysis presented in the following subsection.
((a)) 2.5 m
@@ Appendix B Tree growth evaluation
((c)) 10 m
Figure 11: Tree height dynamics for Mann–Kendall high trend class for three spatial resolutions (2.5 m, 5 m, and 10 m). The curves show median tree height evolution over time with interquartile ranges.
Appendix C Disturbance detection
We evaluate the robustness of our algorithm on the disturbance detection task. Unlike (Fogel et al., 2025), who uses a fixed threshold of 15 m to detect tree loss, we employ an adaptive thresholding approach based on Otsu’s method (Otsu, 1979) to perform binary tree loss classification. Otsu’s method – based on histogram analysis – is commonly used in remote sensing and image analysis to automatically determine a data-driven threshold separating changed and unchanged areas (Kalinicheva et al., 2019). This choice accounts for the fact that canopy changes can occur at more subtle levels than a fixed threshold would capture. For each site, a pixel-wise height difference map ($\Delta H$) is computed between two acquisitions. To reduce noise and limit the influence of single-pixel outliers, a median filter with a $3\times 3$ window is applied. Only negative $\Delta H$ values are considered for threshold computation to obtain disturbance maps.
For postprocessing, as in (Fogel et al., 2025), we applied erosion and dilation with a 3-pixel kernel to regularize the binary masks, and removed connected components smaller than 200 m². This approach enabled robust detection of tree loss while preserving canopy structure and minimizing the influence of noise in the LiDAR data.
Note that tree loss detection is not the primary objective of the proposed algorithm. The disturbance detection workflow presented here is not intended as a final end-to-end solution; in operational settings, more advanced approaches that go beyond simple height differencing are typically required to achieve robust tree loss detection.
In this work, the threshold is computed independently for each site and for each height difference map. This strategy allows the method to be applied consistently across different areas while still producing stable results. In practical applications, some parameters may need to be adapted, such as the spatial extent over which the threshold is estimated.
Nevertheless, we consider this approach more robust than the use of a fixed threshold, as the adaptive Otsu threshold is computed separately for each product and therefore accounts for the specific characteristics of the predicted map. For instance, if a product systematically underestimates height changes compared to the reference data, the Otsu threshold will adjust accordingly.
We evaluate the proposed method on four different study areas: Mouterhouse, Deodatie, Lajoux-Fresse, and Carcans-Hourtins. For each site, an independent ALS-based tree height maps at 1 m resolution are provided by the ONF corresponding to one acquisition year. We match those areas with available LiDAR HD tiles to constitute an image pair for disturbance detection. The four study sites represent different forest compositions and topographic contexts. Carcans-Hourtins, located in the Landes region, is characterized by flat terrain and a maritime pine (Pinus pinaster) dominated stand, with ALS acquisition in October 2020 and LiDAR HD in September 2023. Mouterhouse is also located in a predominantly flat area, although characterized by numerous small forested hills. The ONF ALS data were acquired in February 2019 and the LiDAR HD data in February 2022. The site consists of a mixed deciduous stand dominated by oak, hornbeam, and beech. The Deodatie site is located in a hilly landscape with a mixed forest composition, with ONF ALS data acquired in April 2018 and LiDAR HD in March 2022. Finally, Lajoux-Fresse is situated in a mountainous environment, with ONF ALS acquisition in June 2019 and LiDAR HD in August 2022, and a forest dominated by spruce, Norway spruce, and beech. These sites therefore provide a range of structural conditions, forest types, and topographic settings for evaluating disturbance detection performance.
We evaluate the performance of tree loss detection algorithm based on THREASURE-Net predictions at three different spatial resolutions – 10, 5, and 2.5 m, as well as Open-Canopy (Fogel et al., 2025) at 1.5 m resolution and FORMS-T (Schwartz et al., 2025) at 10 m resolution. For each resolution, the reference ALS height maps are resampled to match the resolution of the prediction in order to ensure a consistent comparison.
The quantitative results are presented in Table 5 and the qualitative results are presented in Figure 12. For each site, we report intersection over union (IoU), precision, recall, and F1-score. Let $TP$ denote true positives, $FP$ false positives, and $FN$ false negatives. Then the metrics are defined as:
$\text{IoU}=\frac{TP}{TP+FP+FN}$
$\text{Precision}=\frac{TP}{TP+FP}$
$\text{Recall}=\frac{TP}{TP+FN}$
$\text{F1-score}=\frac{2\cdot TP}{2\cdot TP+FP+FN}$
As in the previous experiments, the quantitative performance of THREASURE-Net generally decreases as the spatial resolution becomes finer, although the visual results show an increased level of detail. This behavior can be explained by the fact that the reference ALS-derived height difference maps are resampled to match the resolution of the predictions. As a result, at coarser resolutions the reference data contain fewer fine-scale details, which can artificially improve the quantitative metrics.
Overall, the performance of our approach is not systematically better or worse than that of the compared methods. In particular, all algorithms exhibit relatively poor performance at the Mouterhouse site. This may be related to the local topography, which consists of numerous small hills that can introduce additional errors in the height estimates. Nevertheless, at 10 m resolution, our method achieves performance comparable to that of FORMS-T.
In contrast, despite using VHR input data, Open-Canopy did not show the best performance. The visual results indicate that this method tends to underestimate the magnitude of height changes.
Table 5: Tree loss detection performance for for different sites. We report the performance of THREASURE-Net at three different spatial resolutions, as well as Open-Canopy (Fogel et al., 2025) and FORMS-T (Schwartz et al., 2025).
Site (Years) Filename Resolution IoU (%) Recall (%) Precision (%) F1 (%)
Mouterhouse (2022-2019) THREASURE-Net 10 m 30.6 48.2 45.6 46.9
5 m 24.3 30.2 55.8 39.2
2.5 m 18.7 21.8 56.3 31.4
Open-Canopy 1.5 m 20.7 26.5 48.3 34.2
FORMS-T 10 m 33.7 48.1 52.9 50.4
Déodatie (2022-2018) THREASURE-Net 10 m 53.3 76.0 64.1 69.6
5 m 53.2 72.7 66.5 69.4
2.5 m 51.0 72.7 63.0 67.5
Open-Canopy 1.5 m 67.6 79.7 81.6 80.6
FORMS-T 10 m 56.5 85.6 62.4 72.2
Lajoux-Fresse (2022-2019) THREASURE-Net 10 m 55.3 65.6 77.9 71.2
5 m 46.1 62.4 63.8 63.1
2.5 m 40.5 64.7 52.1 57.7
Open-Canopy 1.5 m 51.4 67.9 67.9 67.9
FORMS-T 10 m 53.5 84.8 59.2 69.7
Carcans-Hourtins (2023-2019) THREASURE-Net 10 m 69.8 75.4 90.4 82.2
5 m 72.1 78.2 90.2 83.8
2.5 m 66.1 77.1 82.3 79.6
Open-Canopy 1.5 m 30.2 30.8 93.7 46.4
FORMS-T 10 m 68.7 77.1 86.3 81.5
Mouterhouse Lajoux-Fresse Déodatie Carcans-Hourtins
((a)) SPOT, 2019
((b)) SPOT, 2022
((c)) HDM ALS
((d)) CM ALS
((e)) Ours 10 m
((f)) Ours 10 m
((g)) Ours 5 m
((h)) Ours 5 m
((i)) Ours 2.5 m
((j)) Ours 2.5 m
((k)) Open-Canopy
((l)) Open-Canopy
((m)) FORMS-T
((n)) FORMS-T
((a)) SPOT, 2019
((b)) SPOT, 2022
((c)) HDM ALS
((d)) CM ALS
((e)) Ours 10 m
((f)) Ours 10 m
((g)) Ours 5 m
((h)) Ours 5 m
((i)) Ours 2.5 m
((j)) Ours 2.5 m
((k)) Open-Canopy
((l)) Open-Canopy
((m)) FORMS-T
((n)) FORMS-T
((a)) SPOT, 2018
((b)) SPOT, 2022
((c)) HDM ALS
((d)) CM ALS
((e)) Ours 10 m
((f)) Ours 10 m
((g)) Ours 5 m
((h)) Ours 5 m
((i)) Ours 2.5 m
((j)) Ours 2.5 m
((k)) Open-Canopy
((l)) Open-Canopy
((m)) FORMS-T
((n)) FORMS-T
((a)) SPOT, 2019
((b)) SPOT, 2023
((c)) HDM ALS
((d)) CM ALS
((e)) Ours 10 m
((f)) Ours 10 m
((g)) Ours 5 m
((h)) Ours 5 m
((i)) Ours 2.5 m
((j)) Ours 2.5 m
((k)) Open-Canopy
((l)) Open-Canopy
((m)) FORMS-T
((n)) FORMS-T
Figure 12: Change detection results across the four sites. For each row, the first two panels show SPOT-6/7 images for the change year pairs (a)-(b). Reference ALS-derived maps are shown as height difference maps (HDM) (c) and binary change maps (CM) (d) at 2.5 m resolution. Predictions from THREASURE-Net at three different resolutions are shown in pairs: HDM and CM for each resolution (e)-(f), (g)-(h), (i)-(j) for 10, 5 and 2.5 m respectively . Open-Canopy (Fogel et al., 2025) is shown as HDM and CM (k)-(l), and FORMS-T (Schwartz et al., 2025) as HDM and CM (m)-(n). The final panel in each row shows the corresponding legend.